# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

### v2

@@ 3. One nonparabolic end for minimal hypersurface
Let $(X^{n+1},g)$ be a complete manifold with bounded geometry and $\operatorname{biRic}_{X}\geqslant 0$. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete two-side weakly stable minimal immersion. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Then $M\backslash K$ admits at most one nonparabolic unbounded component. In particular, $M$ has at most one nonparabolic end.
We establish some preliminary results which are crucial in the proof of Theorem 3.1. In the sequel, we assume that the ambient manifold $(X^{n+1},g)$ always has bounded geometry.
3.1. Schoen–Yau inequality
The following inequality of Schoen and Yau is well known in nonnegative sectional curvature. Note that in dimension 4, Chodosh, Li and Stryker [chodosh-li-stryker] generalized it under assumption $\operatorname{Ric}^{X}_{2}\geqslant 0$. They also mentioned that in higher dimension $n+1$, $\operatorname{Ric}^{X}_{n-1}\geq 0$ plays the same role. The same proof works under the assumption $\operatorname{biRic}_{X}\geqslant 0$; see also [Cheng-Cheung-Zhou-CMC][Lemma 3.2] for the generalized inequality on CMC hypersurfaces. Note that $\operatorname{Ric}^{X}_{n-1}\geq 0$ implies $\operatorname{biRic}_{X}\geq 0.$ It would be interesting to understand more about these two assumptions.
@@ 3.1. Schoen–Yau inequality
Let $(X^{n+1},g)$ be a complete manifold with $\operatorname{biRic}_{X}\geqslant 0$. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete two-side weakly stable minimal immersion. Let $u$ be a harmonic function on $M$. Then
(3.1) $\frac{1}{n}\int_{M}\phi^{2}|A_{M}|^{2}|\nabla u|^{2}+\frac{1}{n-1}\int_{M}\phi^{2}|\nabla|\nabla u||^{2}\leqslant\int_{M}|\nabla\phi|^{2}|\nabla u|^{2}$
for any compactly supported, non-negative function $\phi\in W^{1,2}(M)$ satisfying
$\int_{M}\phi|\nabla u|=0.$
3.2. Infinite volume
The purpose of this subsection is to recall the following result which is first proved for complete orientable minimal hypersurface in $\mathbb{R}^{n+1}$ [Cao-Shen-Zhu-infinitevolume].
@@ 3.2. Infinite volume
3.3. One nonparabolic end
For contradiction, suppose that $\{E_{k}\}$ and $\{\widetilde{E}_{k}\}$ are distinct nonparabolic ends of $M$ adapted to a point $p\in M$ with length scale $L$. Since the ends are distinct, there is a $k_{0}\in\mathbb{N}$ such that $E_{k_{0}}\cap\widetilde{E}_{k_{0}}=\emptyset$.
To rule out two nonparabolic ends, we produce a non-constant harmonic function; see [chodosh-li-stryker][Lemma 4.3] for detailed proof.
@@ 3.3. One nonparabolic end
We now turn to the proof of Theorem 3.1.
Proof of Theorem 3.1.
The proof follows from [Cheng-Cheung-Zhou-CMC][Theorem 3.1]. We sketch it here.
Suppose $E$ and $\widetilde{E}$ are nonparabolic unbounded components of $M\backslash K$. By Lemma 3.4, there exists a nonconstant harmonic function $u$ with finite Dirichlet energy. There exists a point $p\in M$ such that
$0<\int_{B_{1}(p)}|\nabla u|^{2}\leq\int_{B_{R}(p)}|\nabla u|^{2}\leqslant{C}\int_{\partial B_{R}(p)}|\nabla u|.$
Applying the co-area formula yields
$\int_{B_{R}(p)}|\nabla u|\geq C_{1}(R-1),$
which implies that $\int_{M}|\nabla u|=\infty$.
Let $\rho_{0}$ be a smoothing of $d_{M}(p,\cdot)$ satisfying $|\nabla\rho_{i}|<2$. We define
$\phi_{1}(R_{0},R):=\left\{\begin{array}[]{ll}1&x\in\overline{B}_{R_{0}}(p)\\ \frac{R_{0}+R-\rho_{0}(x)}{R}&x\in\overline{B}_{R_{0}+R}(p)\backslash B_{R_{0}}(p)\\ 0&x\in M\backslash B_{R_{0}+R}(p),\end{array}\right.$
and
$\phi_{2}(R_{0},R):=\left\{\begin{array}[]{ll}0&x\in\overline{B}_{R_{0}+R}(p)\\ \frac{R_{0}+R-\rho_{0}(x)}{R}&x\in\overline{B}_{R_{0}+2R}(p)\backslash B_{R_{0}+R}(p)\\ -1&x\in B_{R_{0}+2R+R_{1}}(p)\backslash B_{R_{0}+2R}(p)\\ \frac{\rho_{0}(x)-(R_{0}+3R+R_{1})}{R}&x\in\overline{B}_{R_{0}+3R+R_{1}}(p)\backslash B_{R_{0}+2R+R_{1}}(p)\\ 0&x\in M\backslash B_{R_{0}+3R+R_{1}}(p).\end{array}\right.$
Consider $\phi(t,R_{0},R)=\phi_{1}(R_{0},R)+t\phi_{2}(R_{0},R)$. Observe that choosing $R_{1}$ sufficiently large, we have
$\int_{M}\phi(0,R_{0},R)>0,\quad\int_{M}\phi(1,R_{0},R)<0.$
Therefore, there exists some $t_{0}\in(0,1)$ depending on $R_{0}$ and $R$ such that
$\int_{M}\phi(t_{0},R_{0},R)|\nabla u|=0.$
Applying Lemma 3.2, direct calculation yields
$\begin{split}&\frac{1}{n}\int_{B_{R_{0}}(p)}|A_{M}|^{2}|\nabla u|^{2}+\frac{1}{n-1}\int_{B_{R_{0}}(p)}|\nabla|\nabla u||^{2}\\ &\leqslant\int_{M}|\nabla\phi_{1}|^{2}|\nabla u|^{2}+t^{2}_{0}\int_{M}|\nabla\phi_{2}|^{2}|\nabla u|^{2}\\ &\leqslant\frac{4}{R^{2}}\int_{M}|\nabla u|^{2},\end{split}$
where we use $\langle\nabla\phi_{1},\nabla\phi_{2}\rangle=2$ in the first inequality. Letting $R\to\infty$, we conclude that
$|A_{M}|^{2}|\nabla u|^{2}\equiv 0,\quad|\nabla|\nabla u||^{2}\equiv 0.$
Note that $|\nabla u|$ is constant and $u$ is non-constant. We have that $|\nabla u|\neq 0$ everywhere. Theorem 3.3 implies that
$\int_{\Sigma}|\nabla u|^{2}=\operatorname{Vol}(M)|\nabla u|^{2}=\infty,$
which contradicts with the finiteness of Dirichlet energy. ∎
4. CMC hypersurfaces in $5$-manifolds with positive $k\operatorname{-triRic}$
Combining Lemma 2.3, Lemma 2.4 and Theorem 3.3, we can claim that $M^{4}$ is actually stable by [Barbosa-Berard-twisted-eigenvalue][Theorem 3.5]. Theorefore, without loss of generality, we can assume that $M^{4}$ is simply connected by looking at its universal cover. The parabolic ends of $M$ do not affect the construction of cut-off function on $M$ as we will see in the main proof (see also [chodosh-li-stryker][Remark 5.1]), thus we only need to deal with the nonparabolic end.
In this section, we adapt the argument from [chodosh-li-stryker] to prove a diameter bound that uses the theory of so-called balanced warped $\mu$-bubbles for stable minimal hypersurfaces; see also [Gromov19, chodosh-li-bubble] for relevant statements.
@@ 4.1. Balanced $\mu$-bubbles
We recall some preliminary results on existence and stability for balanced warped $\mu$-bubbles. Given a Riemannian $n$-manifold $(M^{n},g)$ with boundary, $n\leqslant 7$, we assume that a choice of labeling the components of $\partial M$ is $\partial M=\partial_{-}M\cup\partial_{+}M$ where neither of them are empty. For a fixed smooth function $u>0$ on $M$ and a smooth function on $\mathring{M}$ with $h\to\pm\infty$ on ±$\partial_{\pm}M$, consider the following balanced warped area functional, $k\in\mathbb{N}$.
@@
for all Caccioppoli sets $\Omega$ with $\Omega\Delta\Omega_{0}\Subset\mathring{M}$ and positive $k$, where the reference set $\Omega_{0}$ with smooth boundary satisfies
@@
under the conformal change $\widetilde{g}=u^{\frac{2k}{n-1}}g$. The critical point of $\widetilde{\mathcal{A}}$ is a hypersurface with the mean curvature $\tilde{H}=hu^{-\frac{k}{n-1}}$. This is equivalent to the following Lemma 4.3. We find that it is more convenient to work with $\mathcal{A}_{k}(\Omega)$ directly instead of $\widetilde{\mathcal{A}}(\Omega)$ under the conformal change.
@@ 4.1.2. Stability
We list the first and second variation for a balanced warped $\mu$-bubble without proof; see [chodosh-li-bubble] for detailed calculation.
@@
$\begin{split}\frac{d^{2}}{dt^{2}}\bigg{|}_{t=0}\mathcal{A}_{k}(\Omega)&=\int_{\Sigma}|\nabla^{\Sigma}\psi|^{2}u^{k}-\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)\right)\psi^{2}u^{k}\\ &+\int_{M}k(k-1)u^{k-2}\left(\nabla_{\nu}u\right)^{2}\psi^{2}\\ &+\int_{\Sigma}ku^{k-1}\left(\Delta_{M}u-\Delta_{\Sigma}u\right)-\left(\nabla_{\nu}(u^{k}h)\right)\psi^{2}.\end{split}$
4.2. Diameter bound
We now turn to the diameter bound. We will need the following proposition.
@@ 4.2. Diameter bound
Let $A=(a_{ij})$ be a $n\times n$ symmetric matrix and denote $H=tr(A)$, we have
@@
for any $k\geq\frac{3}{2}.$ Equality holds if and only if
$a_{11}=a_{nn}=-\frac{2(n-2k-2)}{2nk-3n+6}H,\ \ a_{ii}=\frac{2k-1}{2nk-3n+6}H,\ \text{for}\ i\neq 1,n$
and $a_{ij}=0$ for any $i\neq j.$
@@
Since
$|A|^{2}\geq(a_{11}^{2}+a_{nn}^{2})+\sum_{i=2}^{n-1}a_{ii}^{2}+2(\sum_{i=2}^{n}a_{1i}^{2}+\sum_{i=2}^{n-1}a_{in}^{2})$
By Cauchy–Schwarz inequality, we have
@@
$\geq(k-\frac{3}{2})(a_{11}^{2}+a_{nn}^{2})+\frac{k}{n-2}H^{2}-\frac{2kH}{n-2}(a_{11}+a_{nn})+\frac{k}{n-2}(a_{11}+a_{nn})^{2}$
$+H(a_{11}+a_{nn})$
$\geq(\frac{k}{2}-\frac{3}{4}+\frac{k}{n-2})(a_{11}+a_{nn})^{2}+(1-\frac{2k}{n-2})H(a_{11}+a_{nn})+\frac{k}{n-2}H^{2}.$
The conclusion follows from calculating the minimum. When the equality holds, $a_{11}=a_{nn}$ and $a_{ii}$ are all equal for $i=2,\cdots,n-1$. ∎
We first prove a diameter bound for the warped $\mu$-bubble.
@@
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold. Let $M^{4}\hookrightarrow(X^{5},g)$ be a (strongly) stable two-sided CMC hypersurface with compact boundary $\partial M$ and of the mean curvature $H_{M}$. Suppose
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. There are universal constants $L>0$ and $c>0$ such that if there is a $p\in M$ with $d_{M}(p,\partial M)>\frac{L}{2}$, then in $M$ there is a relatively open set $\Omega\subset B^{M}_{L/2}(\partial M)$ and a smooth closed $3$-dimensional Riemannian manifold $\Sigma^{3}$ (possibly disconnected) such that $\partial\Omega=\Sigma\cup\partial M$ and each component of $\Sigma$ has diameter at most $c$.
@@
Let $\varphi_{0}$ be a smoothing of the function $d(\partial M,\cdot)$ so that $|\nabla\varphi_{0}|\leqslant 2$ and $\varphi_{0}\big{|}_{\partial M}\equiv 0$. Let $\epsilon\in(0,\frac{1}{2})$ so that $\epsilon$ and $\frac{4}{\sqrt{2\varepsilon_{0}}}\pi+2\epsilon$ are regular values of $\varphi_{0}$.
@@
$\varphi=\frac{\varphi_{0}-\epsilon}{\frac{4}{\sqrt{2\varepsilon_{0}}}+\frac{\epsilon}{\pi}}-\frac{\pi}{2}.$
$\Omega_{1}=\{x\in M:-\frac{\pi}{2}<\varphi<\frac{\pi}{2}\}$, and $\Omega_{0}=\{x\in\Sigma^{{}^{\prime}}:-\frac{\pi}{2}<\varphi\leqslant 0\}$. Clearly $|\operatorname{Lip}(\varphi)|\leqslant\frac{\sqrt{2\varepsilon_{0}}}{2}$, and we define
$h=-\sqrt{2\varepsilon_{0}}\tan(\varphi).$
@@
$\nabla h=-\sqrt{2\varepsilon_{0}}\left(1+\tan^{2}(\varphi)\right)\nabla\varphi=-\sqrt{2\varepsilon_{0}}\left(1+\frac{h^{2}}{2\varepsilon_{0}}\right)\nabla\varphi,$

### v3

@@ 3.1. Splitting theorem and finite ends
Let $(X^{n+1},g)$ ($n\leq 5$) be a complete manifold and
$\alpha\text{-}\operatorname{biRic}_{X}+\frac{4-(n-1)\alpha^{2}}{4((1-\alpha)n+\alpha)}H^{2}\geqslant 0$
for $\alpha\in(\frac{n-1}{4},\frac{n}{n-1})$. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete stable (two-sided if it is minimal) CMC immersion with the mean curvature $H$. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Then $M\backslash K$ admits at most two unbounded components. In particular, $M$ has at most one non-parabolic end.
In particular, in order to set $\alpha=1$, the dimension $n$ must be smaller than $5$. The following result in minimal case was first shown in [antonelli-pozzetta-xu, Corollary 1.2].
Theorem 3.2.
Let $(X^{n+1},g)$ ($n\leq 4$) be a complete manifold with
$\operatorname{biRic}_{X}+\frac{5-n}{4}H^{2}\geqslant 0.$
Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete stable (two-sided if it is minimal) CMC immersion with the mean curvature $H$. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Then $M\backslash K$ admits at most two unbounded components. In particular, $M$ has at most one non-parabolic end.
Therefore, under the assumptions of our main theorems in Section 5, it suffices to consider the case where $M^{n}$ has only one end. We will demonstrate in Section 5 that this unique end is parabolic.
3.2. Finite nonparabolic ends of minimal hypersurfaces in general dimensions
In this subsection, following [Li-Wang-nonnegatively-curved-manifold] we show that in a manifold $(X^{n+1},g)$ with certain conditions, the number of nonparabolic ends of a stable CMC hypersurface $(M^{n},g)$ is at most one. This fact is known to experts. We recall some details of its proof.
Theorem 3.3.
Let $(X^{n+1},g)$ be a complete manifold and $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete stable (two-sided if it is minimal) CMC immersion with the mean curvature $H$. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Assume that
  • (1)
    either $\operatorname{Ric}^{X}_{n-1}\geq 0$ and $H=0$,
  • (2)
    or $\operatorname{biRic}_{X}+\frac{5-n}{4}H^{2}\geqslant 0$ and $X$ has bounded geometry.
Then $M\backslash K$ admits at most one nonparabolic unbounded component. In particular, $M$ has at most one nonparabolic end.
There is no dimension restriction, but only $n\geq 5$ is relevant for our purposes, as the splitting result in lower dimensions has already been established in the previous subsection. Theorem 3.3 was proved under the assumption $(1)$, so we will focus on assumption $(2)$.
The following inequality of Schoen and Yau is well-known for minimal hypersurface under nonnegative sectional curvature. Note that in dimension 4, Chodosh, Li and Stryker [chodosh-li-stryker] generalized it under weaker assumption $\operatorname{Ric}^{X}_{2}\geq 0$. They also mentioned that in higher dimension $n+1$, $\operatorname{Ric}^{X}_{n-1}\geq 0$ plays an analogous role. To the best of the authors’ knowledge, the most general form of this inequality was established in [Cheng-Cheung-Zhou-CMC][Lemma 3.2]. It is worth noting that $\operatorname{Ric}^{X}_{n-1}\geq 0$ implies $\operatorname{biRic}_{X}\geq 0.$
@@ 3.2. Finite nonparabolic ends of minimal hypersurfaces in general dimensions
[Cheng-Cheung-Zhou-CMC, Lemma 3.2] Let $(X^{n+1},g)$ be a complete manifold. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete stable (two-sided if it is minimal) CMC immersion with the mean curvature $H$. Let $u$ be a harmonic function on $M$. Then
(3.1) $\int_{M}(\operatorname{biRic}_{X}(e,\nu)+\frac{5-n}{4}H^{2})\phi^{2}|\nabla u|^{2}+\frac{1}{n-1}\int_{M}\phi^{2}|\nabla|\nabla u||^{2}\leqslant\int_{M}|\nabla\phi|^{2}|\nabla u|^{2}$
for any compactly supported, non-negative function $\phi\in W^{1,2}(M)$ where $e=\nabla u/|\nabla u|.$
Recall the following result which is first proved for complete orientable minimal hypersurfaces in $\mathbb{R}^{n+1}$ [Cao-Shen-Zhu-infinitevolume].
@@
For contradiction, suppose that $\{E_{k}\}$ and $\{\widetilde{E}_{k}\}$ are distinct nonparabolic ends of $M$ adapted to a point $p\in M$ with length scale $L$. We then produce a non-constant harmonic function; see [chodosh-li-stryker][Lemma 4.3] for detailed proof.
@@
By Lemmas 3.4 and 3.6, choosing a standard cutoff function based on the distance function, we can deduce from (3.1) that $\operatorname{biRic}_{X}(e,\nu)+\frac{5-n}{4}H^{2}\equiv 0$ and $|\nabla u|$ is a nonzero constant. This contradicts the finite energy condition of $u$ and Theorem 3.5. Therefore, $M$ has at most one nonparabolic end.
4. CMC hypersurfaces in $5$ or $6$-manifolds
Without loss of generality, we can assume that $M^{n}$ is simply-connected by looking at its universal cover. The parabolic ends of $M$ do not affect the construction of cut-off function on $M$ as we will see in the main proof (see also [chodosh-li-stryker][Remark 5.1]), thus we only need to deal with one nonparabolic end (it possibly does not exist).
In this section, we adapt the argument from [chodosh-li-stryker] to prove a volume bound that uses the theory of so-called balanced warped $\mu$-bubbles for stable CMC hypersurfaces; see also [Gromov19, chodosh-li-bubble] for relevant statements.
@@ 4.1. Balanced $\mu$-bubbles
We recall some preliminary results on existence and stability for balanced warped $\mu$-bubbles. Given a Riemannian $n$-manifold $(M^{n},g)$ with boundary, $n\leqslant 7$, we assume that a choice of labeling the components of $\partial M$ is $\partial M=\partial_{-}M\cup\partial_{+}M$ where neither of them are empty. For a fixed smooth function $u>0$ on $M$ and a smooth function on $\mathring{M}$ with $h\to\pm\infty$ on ±$\partial_{\pm}M$, consider the following balanced warped area functional:
@@
for all Caccioppoli sets $\Omega$ with $\Omega\Delta\Omega_{0}\Subset\mathring{M}$, where the reference set $\Omega_{0}$ with smooth boundary satisfies
@@
under the conformal change $\widetilde{g}=u^{\frac{2k}{n-1}}g$. The critical point of $\widetilde{\mathcal{A}}$ is a hypersurface with the mean curvature $\tilde{H}=hu^{-\frac{k}{n-1}}$. This is equivalent to the statement in Lemma 4.3. We find that it is more convenient to work with $\mathcal{A}_{k}(\Omega)$ directly instead of $\widetilde{\mathcal{A}}(\Omega)$ under the conformal change.
@@ 4.1.2. Stability
We list the first and second variation for a balanced warped $\mu$-bubble without proof; see [chodosh-li-bubble] for detailed calculations.
@@
$\begin{split}\frac{d^{2}}{dt^{2}}\bigg{|}_{t=0}\mathcal{A}_{k}(\Omega)&=\int_{\Sigma}|\nabla^{\Sigma}\psi|^{2}u^{k}-\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)\right)\psi^{2}u^{k}\\ &+\int_{\Sigma}k(k-1)u^{k-2}\left(\nabla_{\nu}u\right)^{2}\psi^{2}\\ &+\int_{\Sigma}ku^{k-1}\left(\Delta_{M}u-\Delta_{\Sigma}u\right)\psi^{2}-\left(\nabla_{\nu}(u^{k}h)\right)\psi^{2}.\end{split}$
4.2. Volume bound for bubbles
In this subsection, we are going to estimate the volume of $\mu$-bubbles constructed in stable two-sided CMC hypersurfaces. We first prove the following two technical propositions.
@@ 4.2. Volume bound for bubbles
Let $A=(a_{ij})$ be an $n\times n$ symmetric matrix with $n\geq 2$, and let $H=tr(A)$. Then, the following holds:
@@
for any
(4.1) $k\geq\max\left\{1,\frac{3(n-2)}{2n}\right\}.$
Equality holds if and only if
$\begin{split}a_{11}&=a_{nn}=-\frac{2(n-2k-2)}{2nk-3n+6}H,\\ a_{ii}&=\frac{2k-1}{2nk-3n+6}H,\quad i\neq 1,n,\\ a_{ij}&=0,\quad~{}~{}\mbox{otherwise}~{}~{}.\end{split}$
@@
Combining Cauchy–Schwarz inequality and the fact
$|A|^{2}\geq(a_{11}^{2}+a_{nn}^{2})+\sum_{i=2}^{n-1}a_{ii}^{2}+2(\sum_{i=2}^{n}a_{1i}^{2}+\sum_{i=2}^{n-1}a_{in}^{2}),$
we have
@@
$\geq$ $\left(\frac{k}{2}-\frac{3}{4}\right)(a_{11}+a_{nn})^{2}+\frac{k}{n-2}H^{2}-\frac{2kH}{n-2}(a_{11}+a_{nn})$
$+\frac{k}{n-2}(a_{11}+a_{nn})^{2}+H(a_{11}+a_{nn})$
$\geq$ $\left(\frac{k}{2}-\frac{3}{4}+\frac{k}{n-2}\right)(a_{11}+a_{nn})^{2}+\left(1-\frac{2k}{n-2}\right)H(a_{11}+a_{nn})+\frac{k}{n-2}H^{2}.$
The conclusion now follows from elementary calculations. When equality holds, we have $a_{11}=a_{nn}$ and the diagonal entries $a_{ii}$ are all equal for $i=2,\cdots,n-1$. ∎
Notice that $\frac{3(n-2)}{2n}<1$ when $n\leq 5$. Thus, when applying Proposition 4.5, we can merely assume $k\geq 1$ when $n\leq 5.$ We have a more general result as follows.
Proposition 4.6.
Let $A=(a_{ij})$ be an $n\times n$ symmetric matrix and denote $H=\operatorname{tr}(A)$. For any $k\geq 1/2$ and $\alpha\leq 2k$ satisfying
(4.2) $\begin{split}&k-\frac{(n-2)\alpha^{2}}{4k}>0,\\ \quad k^{2}-&\frac{1}{4}-\frac{n-2}{4}\left(1+\alpha^{2}-\frac{\alpha}{k}\right)>0,\end{split}$
there exists a positive constant $\delta=\delta(n,k,\alpha)$ such that
$I:=k|A|^{2}-\alpha a_{11}a_{nn}-\alpha\sum_{i=1}^{n-1}a_{1i}^{2}-\sum_{i=1}^{n}a_{in}^{2}+H(\alpha a_{11}+a_{nn})\geq\frac{\delta}{n}H^{2}.$
Proof.
We consider the matrix
(4.3) $P=\begin{pmatrix}k&\frac{1}{2}&\frac{1}{2}&\cdots&\frac{1}{2}\\ \frac{1}{2}&k&\frac{\alpha}{2}&\cdots&\frac{\alpha}{2}\\ \frac{1}{2}&\frac{\alpha}{2}&k&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{2}&\frac{\alpha}{2}&0&\cdots&k\end{pmatrix}_{n\times n}.$
By elementary row and column operations, it is easy to see that the matrix $P$ is congruent to
$\begin{pmatrix}k-\frac{n-2)}{4k}&\frac{1}{2}\left(1-\frac{(n-2)\alpha}{2k}\right)&0&\cdots&0\\ \frac{1}{2}\left(1-\frac{(n-2)\alpha}{2k}\right)&k-\frac{(n-2)\alpha^{2}}{4k}&0&\cdots&0\\ 0&0&k&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&k\end{pmatrix}_{n\times n}.$
Therefore, the positive definiteness of $P$ is equivalent to the positive definiteness of $Q$:
$Q=\begin{pmatrix}k-\frac{n-2}{4k}&\frac{1}{2}\left(1-\frac{(n-2)\alpha}{2k}\right)\\ \frac{1}{2}\left(1-\frac{(n-2)\alpha}{2k}\right)&k-\frac{(n-2)\alpha^{2}}{4k}\end{pmatrix}$
which requires
(4.4) $\begin{split}k&-\frac{(n-2)\alpha^{2}}{4k}>0\\ \det Q=k^{2}-\frac{1}{4}-&\frac{n-2}{4}\left(1+\alpha^{2}-\frac{\alpha}{k}\right)>0.\end{split}$
Then we set
$x=\begin{pmatrix}a_{nn}\\ a_{11}\\ a_{22}\\ \vdots\\ a_{n-1,n-1}\\ \end{pmatrix}.$
By assumptions (4.2) and Cauchy–Schwarz inequality, we have
$I\geq x^{T}Px\geq\delta|x|^{2}\geq\frac{\delta}{n}H^{2}$
where $\delta$ is the smallest eigenvalue of $P$.
We now work on the volume estimate of the warped $\mu$-bubble.
@@
Let $(X^{n+1},g)$ $(4\leq n\leq 5)$ be a complete $(n+1)$-dimensional Riemannian manifold, and let $M^{n}\hookrightarrow(X^{n+1},g)$ be a stable two-sided CMC hypersurface with compact boundary $\partial M$ and the mean curvature $H_{M}$. Suppose that there exists $k,\alpha$ satisfying (4.1) or (4.2). Moreover, assume that $k,\alpha$ satisfy
(4.5) $\frac{4}{(4-k)\alpha}\leq\frac{n-2}{n-3},\quad G\geq 0,$
where the matrix $G$ is given in (4.10).
Furthermore, if we have
(4.6) $\lambda_{k,\alpha}+C(k,n)H_{M}^{2}\geqslant 2\varepsilon_{0}>0,$
where $\lambda_{k,\alpha}$ is the infimum of the $(k,\alpha)$-triRicci curvature on $X^{n+1}$ and the constant $C(k,n)$ is given by
(4.7) $C(k,n):=\left\{\begin{array}[]{cc}\frac{2k^{2}+k-n+2}{2nk-3n+6},&\alpha=1\\ \delta(k,n,\alpha),&~{}~{}\mbox{otherwise}~{}~{},\end{array}\right.$
then there exist universal constants $L>0$ and $c>0$ such that if $M$ contains a point $p$ satisfying $d_{M}(p,\partial M)>\frac{L}{2}$, then there exists a relatively open set $\Omega\subset B^{M}_{L/2}(\partial M)$ and a smooth closed $(n-1)$-dimensional Riemannian manifold $\Sigma^{n-1}$ (possibly disconnected) such that $\partial\Omega=\Sigma\cup\partial M$ and the volume of each component, still denoted by $\Sigma$, is bounded by $c$.
Remark 4.8.
In this remark, we discuss the admissible choice of $k$ and $\alpha$ that meets several restrictions in the proof of Lemma 4.7.
  • (1)
    When $n=4$, we may choose $k=1$, $\alpha=1$. It satisfies the assumption (4.1). For the matrix $G$ defined in (4.10), we can take $\beta=1/3$ such that
    $\frac{1}{k}-\beta=\frac{2}{3}>0,\quad 1-\alpha\frac{n-2}{n-1}=\frac{1}{3}>0,\quad\det G=\frac{1}{36}>0,$
    i.e., $G$ is positive definite. Note that $\frac{4}{(4-k)\alpha}\leqslant 2=\frac{n-2}{n-3}$ which verifies (4.5). By [antonelli-xu, Theorem 1], there exists a universal constant $c>0$ such that
    $\operatorname{Vol}(\Sigma)\leqslant c.$
    On the other hand, condition (4.6) becomes
    $C_{3}+\frac{1}{2}H_{M}^{2}\geq 2\varepsilon_{0}$
    where $C_{3}$ is the $3$-intermediate curvature.
  • (2)
    When $n=5$, the choice of $k$ and $\alpha$ becomes very delicate. First, it is not hard to check that when $\alpha=1$, there does not exist $k\geq 1$ satisfying (4.5) and (4.1). Moreover, when $\alpha$ is slightly bigger than $1$, condition (4.2):
    $\quad k^{2}+\frac{3\alpha}{4k}>\frac{3}{4}\alpha^{2}+1>\frac{3}{4}\alpha+1$
    requires $k>1$ but the condition $\det G>0$:
    $\frac{1}{k}>\frac{\beta\left((1-\frac{3\alpha}{4})(\frac{3\alpha}{16}-\frac{3}{4})-\frac{3\alpha^{2}}{64}\right)+\frac{1}{4}(1-\frac{3\alpha}{4})}{-\frac{3}{16}\alpha^{2}+\left(\frac{3}{16}-\frac{3}{4}(1-\beta)\right)\alpha-\frac{3}{16}+1-\beta}$
    requires $k<1$. In all, we may choose $\alpha=0.99$, $k=1.001$, and $\beta=0.001$. Direct calculations yield
    $\begin{split}k-\frac{3}{4k}\alpha^{2}&=0.2667,\\ k^{2}-\frac{1}{4}-\frac{3}{4}\left(1+\alpha^{2}-\frac{\alpha}{k}\right)&=0.0087,\end{split}$
    which verifies assumptions (4.2). Besides, we have
    $\frac{1}{k}-\beta=0.998>0,\quad 1-\alpha\frac{3}{4}=0.2575>0,\quad\det G=0.0017.$
    So $G$ is positive definite. Moreover, $\frac{4}{(4-k)\alpha}=1.3472\leq\frac{3}{2}=\frac{n-2}{n-3}$ satisfies (4.5), we have by [antonelli-xu, Theorem 1] that
    $\operatorname{Vol}(\Sigma)\leqslant c.$
@@
Let $\varphi_{0}$ be a smoothing of the function $d(\partial M,\cdot)$ so that $|\nabla\varphi_{0}|\leqslant 2$ and $\varphi_{0}\big{|}_{\partial M}\equiv 0$. Let $\epsilon\in(0,\frac{1}{2})$ so that $\epsilon$ and $\frac{2}{\sqrt{\beta\varepsilon_{0}}}\pi+2\epsilon$ are regular values of $\varphi_{0}$, where the positive constant $\beta$ will be determined later.
@@
$\varphi=\frac{\varphi_{0}-\epsilon}{\frac{2}{\sqrt{\beta\varepsilon_{0}}}+\frac{\epsilon}{\pi}}-\frac{\pi}{2}.$
$\Omega_{1}=\{x\in M:-\frac{\pi}{2}<\varphi<\frac{\pi}{2}\}$, and $\Omega_{0}=\{x\in\Sigma^{{}^{\prime}}:-\frac{\pi}{2}<\varphi\leqslant 0\}$. Clearly $|\operatorname{Lip}(\varphi)|\leqslant\sqrt{\beta\varepsilon_{0}}$, and we define
$h=-\sqrt{\frac{\varepsilon_{0}}{\beta}}\tan(\varphi).$
@@
$\nabla h=-\sqrt{\frac{\varepsilon_{0}}{\beta}}\left(1+\tan^{2}(\varphi)\right)\nabla\varphi=-\sqrt{\frac{\varepsilon_{0}}{\beta}}\left(1+\frac{\beta h^{2}}{\varepsilon_{0}}\right)\nabla\varphi,$